OpenStudy (anonymous):

If f(x) = x2 − 2x + 9 and g(x) = 8 − x, what is (f o g)(−4)? Select one of the options below as your answer: A. 111 B. 144 C. 120 D. 129

OpenStudy (shamil98):

Evaluate g(x) at -4, then input that into f(x)

OpenStudy (anonymous):

g(-4) = 12 Now write f(12). (12)^2-2(12) + 9 = 129

OpenStudy (anonymous):

It's called composition. You could also write it as: f(g(x)) = (8-x)^2 -2(8-x) + 9 Now sub in -4 for x. f(g(x)) = 129